OpenStudy (anonymous):

PLEASE HELP!!!!! The midpoint of uv is (5 -11). the coordinates of one endpoint are u(3,5). find the coordinates of endpoint v. please explain!!

OpenStudy (amistre64):

what is your concept of a midpoint?

OpenStudy (amistre64):

midpoint is halfway between 2 other points 5 is halfway between 3 and what value? -11 is halfway between 5 and what value? if you know how to take an average, then 5 and -11 are your averages for x and y

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

halfway should be simple enough to determine ... if you start at 3, and walk to 5, how far have you walked?

OpenStudy (amistre64):

 halfway is the same distance from the end points ..... if you know how far you have come, you know how far you need to go to get to the other side.

OpenStudy (anonymous):

you can always use this :) Since we know that mid point formula is: \[M=(\frac{ x1 + x2 }{ 2 },\frac{ y1+y2 }{ 2 }) \] now we know that your end point is U (3,5), can say that this one is x1,y1 or x2,y2 to you doesn't matter, just choose one and find the other. I use the midpoint as the results since we're trying to find the other point. Cause it like to find x when you have x+y=z and y =7, z=10. ( if you still don't understand, please feel free to tag me or ask again so I could explain more properly) so, it should look like this: \[( \frac{ 3+x2 }{ 2 }) = 5 \] 3+x2=10 x2=7 \[(\frac{ 5+y2 }{ 2 }) = -11\] 5+y2=-22 y2=-27 so the other end point if (7,-22)